# Robot-mediated physical Human-Human Interaction in Neurorehabilitation: a position paper

### v1

@@ Abstract
Neurorehabilitation conventionally relies on the interaction between a patient and a physical therapist. Robotic systems can improve and enrich the physical feedback provided to patients after neurological injury, but they under-utilize the adaptability and clinical expertise of trained therapists. In this position paper, we advocate for a novel approach that integrates the therapist’s clinical expertise and nuanced decision-making with the strength, accuracy, and repeatability of robotics: Robot-mediated physical Human–Human Interaction. This framework, which enables two individuals to physically interact through robotic devices, has been studied across diverse research groups and has recently emerged as a promising link between conventional manual therapy and rehabilitation robotics, harmonizing the strengths of both approaches. This paper presents the rationale of a multidisciplinary team—including engineers, doctors, and physical therapists—for conducting research that utilizes: a unified taxonomy to describe robot-mediated rehabilitation, a framework of interaction based on social psychology, and a technological approach that makes robotic systems seamless facilitators of natural human-human interaction.
@@ I Position
The common trope regarding robots is that they could gradually replace skilled human jobs. The same concern arises in neurorehabilitation (NR). The use of robots in rehabilitation settings offers exercise automation and reduced therapist effort, but these benefits are overshadowed by the limited impact they currently have on improving motor learning and incorporating clinician input [1, 2, 3]. In this position paper, we present a rationale for using robots in NR with the ultimate goal of improving clinical outcomes. To achieve this, we propose leveraging robotics to facilitate therapist-patient physical interaction in a way that preserves the nuanced, skilled care provided by experienced physical therapists. The integration of these novel technologies, which combine the adaptability, flexibility, and social interaction of physical human-human interaction with the physical strength and controllability of robots, is poised to have a transformative impact on clinical care. In NR, this approach can enhance functional training, increase therapy dose, improve the overall quality of care, and help reduce healthcare costs.
Figure 1: Examples of Robot-mediated physical Human-Human Interaction (RHHI) in neurorehabilitation (NR): (A) Two users (including at least one patient) can interact with two robots. The patients can interact with a therapist, another patient, or a family member (Partner Characteristics). Task definition influences the strategies implemented by the two users and their relative performance. (B) Multiple patients can perform interactive tasks by virtually connecting multiple devices; in this example, the patients interact through their wrist joints.
@@ II Background
Millions of people around the world face significant challenges due to sensorimotor impairments associated with neurological conditions such as stroke and spinal cord injury (SCI) [4, 5]. These impairments severely restrict their independence, employment opportunities, and diminish overall quality of life [6, 7], while also imposing a profound economic burden on families and healthcare systems [8, 9, 10].
For these reasons, individuals with sensorimotor impairments routinely engage in functional rehabilitation exercises in clinical settings to regain control of their movements [11]. These exercises range from upper-limb tasks, such as handwriting or grasping objects, to full-body coordination activities, such as sit-to-stand transitions and overground walking. Physical therapists play a crucial role in the rehabilitation process. They collaborate with clinicians in the design of individualized rehabilitation plans and are trained to tailor interventions to each patient’s specific needs, adapting their strategies based on the patient’s condition and progress.
Importantly, therapists are trained to use physical contact to assess and treat their patients. Through tactile feedback, they can evaluate a patient’s abilities, identify areas for improvement, and offer direct interventions by physically guiding their patient’s limbs during exercises. Thus, physical therapists represent a critical source of extrinsic feedback for patients, while also providing implicit proprioceptive input through physical interaction—such as guiding a joint, applying resistance, or supporting a limb—which can complement or even partially compensate for compromised intrinsic feedback due to impaired somatosensory pathways. In doing so, they facilitate the internalization of correct motor patterns [12] and ensure task completion by providing appropriate assistance, which helps avoid frustration, reinforce success, and sustain motivation during high-intensity, repetitive training [13].
@@
Nevertheless, helping patients with whole-body movements or severe impairments places a significant physical burden on therapists. While high-intensity, repetitive training is essential for effective rehabilitation, it further adds to this strain. Tasks like sit-to-stand transfers or ambulation are especially challenging when patients require moderate assistance to maintain balance [14, 15]. This physical strain increases the risk of musculoskeletal injuries [16] and can limit the precision and force therapists are able to apply. In some cases, optimal care requires multiple therapists, compromising therapist availability. This overall complexity also makes accurate measurement difficult and complicates the characterization and repetition of therapeutic procedures [17, 18].
@@
Rehabilitation robots help address these clinical challenges by isolating impaired joint(s) of interest and providing precise, repetitive practice through assistive or resistive forces. They ensure safety via compliant control strategies and built-in force and motion limits, while promoting neuroplasticity through consistent and functional movement practice [19]. Furthermore, they facilitate the assessment of patient performance and allow objective quantification of progress throughout the rehabilitation process. When combined with real-time visual feedback and gamified elements, rehabilitation robots can further enhance therapy by informing and motivating patients about their performance, thereby encouraging more active participation and engagement, ultimately improving therapeutic outcomes [20, 21]. Nevertheless, the use of rehabilitation robots often entails the loss of physical contact between therapists and patients, removing a critical channel for understanding each other’s intentions and promoting motor skill acquisition. In this case, a less experienced or proactive therapist may retreat into passive supervision and simple parameter tweaks, resulting in an under-utilization of both the robot’s capabilities and their own clinical expertise [1, 22].
@@
This paper highlights a novel framework that integrates therapist expertise with the strength and repeatability of robotics through Robot-mediated physical Human-Human Interaction (RHHI). Enabling two individuals to physically interact through robotic devices, RHHI has emerged as a promising bridge between conventional manual interventions and robotic rehabilitation. We outline the key benefits of this approach, examine its current limitations, and propose directions for future research and development. The paper is structured as follows: Sec. III introduces the salient aspects of RHHI and a taxonomy of this novel framework. Sec. IV presents how this approach can be used in NR to simultaneously leverage the strengths of therapists and robotic devices. Sec. V outlines examples of potential applications of these technologies in the context of conventional rehabilitation practice. Sec. VI discusses challenges associated with the implementation of these technologies. Finally, Sec. VII explores open questions and future research directions.
Figure 2: Examples of RHHI in NR: (A) Therapist and patient interact via two upper-limb exoskeletons. They can assume different roles during the interaction (Interaction Scenario); (B) Robotic devices can be used for learning by demonstration strategies: a therapist demonstrates a movement to the robot which learns how to replicate that movement in autonomy.
@@ III Taxonomy and Overview of RHHI in NR
In RHHI, one or more robotic devices are used to render specific interaction dynamics between individuals [23, 24, 25]. In NR, at least one individual is a patient physically connected to a robot, whose movements are influenced in real time by another user, such as a therapist, another patient, or even a family member. Together, the individuals engage in coordinated movements aimed at promoting motor learning for the patient(s) and facilitating the transfer of these skills to everyday functional activities. RHHI’s versatility supports a wide range of scenarios and experimental designs, making it a flexible and powerful framework for advancing physical human-human interaction in NR. To better structure and analyze these scenarios, the following sections introduce key characterizations of RHHI with specific attention to their impact in NR.
@@ III-A Interaction Configurations
Task-Space interactions, where users exchange forces through the end-effector of a robot (Fig. 1.A and Fig. 3.C). This configuration allows force exchange at the end-effector but does not provide specific postural corrections. Single-joint interactions, the two users interact by moving a single joint (e.g., in Fig. 1.B, the wrist). This interaction is mediated using a single-DoF robots and allows joint-specific training. Multi-joint interactions allows coordination across multiple DoFs (Fig. 2.A and Fig. 3.A). This type of interaction is typically achieved with exoskeletons. Cross-limb interactions, in which the upper body of one user is connected to the lower body of another (Fig. 3.C). Guided interactions, where one person interacts with a robot and another person can influence the robot’s behavior, for example, by teaching a kinematic or force profile while directly interacting with the robot (Fig. 2.B).
RHHI can accommodate various interaction configurations, including bidirectional and unidirectional interaction. The interaction is defined as bidirectional if both users perceive interaction forces due to the interaction between them, promoting embodiment. This configuration can be implemented by either using two robotic devices that both render interaction forces to each user (Fig. 3.A) or using a transparent robotic device that directly transmits forces applied by one user to a second user (Fig. 2.B). Unidirectional interaction refers to a setup in which only one user perceives forces through a robotic device, while the device’s motion is teleoperated or influenced by another user. This can be done with a single device, teleoperated or influenced by another user, or with two similar devices where one device is the leader, and the other is the follower. This interaction can occur in real time or using prior demonstrations and replication (Fig. 3.B).
@@ III-B Interaction Medium
In RHHI, the interaction medium refers to the virtual connection implemented between two users through robotic devices. This connection can be designed to render specific mechanical behaviors, such as a virtual spring-damper model between joint positions, that attracts the partners’ joints toward a shared equilibrium position. This approach enables users to feel each other’s movements through adjustable mechanical impedance levels [24, 26, 23, 27, 28]. A rigid virtual spring constant results in closely coupled motion, where users exhibit nearly identical kinematic profiles [24, 29]. In contrast, a compliant, softer virtual spring allows more independent movement while still facilitating the exchange of forces. An asymmetric interaction medium can also be implemented, such as the case where one user experiences a rigid connection while the other experiences a soft one [30], enabling customized feedback and workload distribution. It is important to note that the resulting behavior is not solely determined by the virtual model, but also by the human biomechanics, such as the stiffness of the involved joints, residual strength, and active participation of each user. Indeed, individuals can modulate their joint stiffness to either resist or follow the forces transmitted through the interaction [31].
Figure 3: Examples of RHHI in NR: A therapist and patient physically interact in different configurations. (A) The two bi-directionally exchange forces through two lower-limb exoskeletons. The interaction medium between the two users is modeled using a virtual spring-damper system characterized by therapist and patient stiffness ($K_{T},K_{P}$) and damping ($D_{T},D_{P}$). (B) The two interact unidirectionally. This means only the patient perceives forces; in this example, the therapist wears IMU sensors to teleoperate the patient’s exoskeleton. Both (A),(B) represent examples of interaction in joint space while (C) represents an example in task space, as well as across different joints (i.e., the therapist’s upper-limb interacting with the patient’s lower-limb).
@@ III-C Partner characteristics
Partner characteristics refer to individual and relational factors that influence how users behave during RHHI. These include the relative skill levels of each user, as well as their social relationship (e.g., strangers, friends, family members, peers, or in a patient-therapist dynamic) [26, 32, 23, 33, 34, 35, 36, 37, 38]. Such characteristics can affect muscular effort, response strategies, task engagement, and the roles that emerge during the interaction. For example, a more skilled or confident partner may naturally take on a guiding role, while a less experienced partner may assume a more passive role.
Another important dimension of the framework is the number of users participating in RHHI. Most studies focus on dyadic interactions, in which two individuals are physically connected through one or two robotic devices. However, multi-user configurations have also been explored with examples in literature up to four users [34]. While these setups present additional challenges in terms of control complexity and task coordination, they offer unique opportunities for social engagement, scalability, and group-based rehabilitation.
@@ III-D Interaction Scenario
The interaction scenario defines the way the task is structured and how roles are assigned between participants [39]. Common scenarios include collaboration, cooperation, and competition. These modes of interaction can significantly shape user behavior, influencing engagement, learning strategies, and the physical dynamics of interaction. In collaborative scenarios, both users share the same task goal and work together without predefined roles [26]. For example, a patient may perform a task jointly with a therapist, another patient, or a family member, sharing the effort required to complete movements that might otherwise be too difficult alone. In cooperative scenarios, users still share a goal, but with distinct roles such as teacher/student [40, 41, 42] or leader/follower [43]. A therapist, for instance, may lead the interaction by providing guidance while the patient adopts a more passive role. In contrast, competitive scenarios involve opposing goals, often implemented through game-based tasks [38]. Competitive strategies can be employed to encourage patients’ active engagement and augment the challenge of a given task.
@@ III-E Task Performance and Motor Learning
Task performance in RHHI quantifies how users perform the assigned motor task while physically connected. Performance is typically measured through metrics such as accuracy, timing, smoothness, or synchronization. It is influenced by the design of the interaction medium, the characteristics of the users, and the selected interaction scenario. Physical feedback exchanged between users has been shown to improve accuracy, smoothness, coordination and overall task efficiency [44, 45, 24, 46].
Motor learning refers to persistent improvements in motor performance due to the RHHI training, after the physical connection is removed. This lasting change represents the core therapeutic effect in NR, reflecting the patient’s ability to retain and apply improved motor strategies independently. This is often assessed by comparing individual performance (i.e., without the physical interaction) before and after training. RHHI has been shown to support motor learning through various mechanisms, including increased motivation and engagement [37, 47, 48], the opportunity to estimate and internalize a partner’s movement strategy [49, 24, 34] and role specialization that emerges during interaction [44, 45]. These benefits make RHHI particularly promising for NR, where one of the users aims to regain lost motor functions through repeated, adaptive practice.

### v2

# Robot-mediated physical Human-Human Interaction in Rehabilitation: a position paper

@@ Abstract
Neurorehabilitation conventionally relies on the interaction between a patient and a physical therapist. Robotic systems can improve and enrich the physical feedback provided to patients after neurological injury, but they under-utilize the adaptability and clinical expertise of trained therapists. In this position paper, we advocate for a novel approach that integrates the therapist’s clinical expertise and nuanced decision-making with the strength, accuracy, and repeatability of robotics: Robot-mediated physical Human–Human Interaction. This framework, which enables two individuals to physically interact through robotic devices, has been studied across diverse research groups and has recently emerged as a promising link between conventional manual therapy and rehabilitation robotics, harmonizing the strengths of both approaches. Although current findings are largely based on pilot studies and conceptual frameworks, integrating therapists’ expertise with the functionalities offered by robotic systems represents a promising direction for improving rehabilitation outcomes. This paper presents the rationale of a multidisciplinary team—including engineers, doctors, and physical therapists—for conducting research that utilizes: a unified taxonomy to describe robot-mediated rehabilitation, a framework of interaction based on social psychology, and a technological approach that makes robotic systems seamless facilitators of natural human-human interaction.
@@ I Position
The common trope regarding robots is that they could gradually replace skilled human jobs. The same concern arises in neurorehabilitation (NR). The use of robots in rehabilitation settings offers exercise automation and reduced therapist effort. However, these benefits are overshadowed by the limited impact they currently have on improving motor learning and incorporating clinician input [celian2021day, tamburella2024quantifying, yamamoto2022effect]. In this position paper, we present a rationale for using robots in NR with the ultimate goal of improving clinical outcomes. To achieve this, we propose leveraging robotics to facilitate therapist-patient physical interaction in a way that preserves the nuanced, skilled care provided by experienced physical therapists. The integration of these novel technologies, which combine the adaptability, flexibility, and social interaction of physical human-human interaction with the physical strength and controllability of robots, is poised to have a transformative impact on clinical care. To achieve this long-term goal, statistically powered clinical studies are necessary to validate the expected improvements in functional training, therapy dose, quality of care, and healthcare cost reduction.
@@ II Background
Millions of people around the world face significant challenges due to sensorimotor impairments associated with neurological conditions such as stroke and spinal cord injury (SCI) [feigin2025world, singh2014global]. These impairments severely restrict their independence, employment opportunities, and diminish overall quality of life [carod2009quality, post2005quality], while also imposing a profound economic burden on families and healthcare systems [diop2021quality, strilciuc2021economic, lucas2023costs].
Individuals with sensorimotor impairments routinely engage in functional rehabilitation exercises in clinical settings to regain control of their movements, guided by physical therapists who play a central role by interpreting patient states, modulating assistance, and ensuring safety [shumway2007motor, higgs2024clinical]. They design and adapt individualized rehabilitation plans in collaboration with clinicians, tailoring interventions to each patient’s condition and progress. Through empathy and affective perception, therapists provide personalized feedback and adjust therapy based on patient readiness, fatigue, and emotional state [wulf2016optimizing, deci2008self].
Importantly, therapists are trained to use physical contact to assess and treat patients. Through haptic feedback, they evaluate motor abilities, identify deficits, and guide limb movements during exercises. Physical contact provides both explicit feedback and implicit proprioceptive input, which can complement or compensate for impaired somatosensation. This interaction facilitates the internalization of correct motor patterns [hornby2020clinical] and ensures task completion and reinforcing success during intensive, repetitive training [marchal2009review].
@@
Helping patients with whole-body movements or severe impairments places a significant physical burden on therapists. While high-intensity, repetitive training is essential for effective rehabilitation, it further adds to this strain [burnfield2013comparative, boukadida2015determinants]. This physical strain increases the risk of musculoskeletal injuries [mccrory2014work] and can negatively affect task execution, e.g., by reducing smoothness and coordination, and may further hinder the therapist’s ability to effectively supervise the rehabilitation process. In some cases, optimal care requires multiple therapists, compromising therapist availability. This overall complexity also makes accurate measurement difficult and complicates the characterization and repetition of therapeutic procedures [sawers2014perspectives, galvez2005measuring].
@@
Rehabilitation robots help address these challenges by providing precise, repetitive practice through assistive or resistive forces [gassert2018rehabilitation]. They facilitate the assessment of patient performance and allow objective quantification of progress throughout the rehabilitation process. However, the use of rehabilitation robots often entails the loss of physical contact between therapists and patients, removing a critical channel for understanding each other’s intentions and promoting motor skill acquisition. In this case, a less experienced therapist may retreat into passive supervision and simple parameter tweaks, resulting in an under-utilization of both the robot’s capabilities and their own clinical expertise [celian2021day, postol2023we].
@@
This paper highlights a novel framework that integrates therapist expertise with the strength and repeatability of robotics through Robot-mediated physical Human-Human Interaction (RHHI). The paper is structured as follows: Sec. III introduces the salient aspects of RHHI and a taxonomy of this novel framework. Sec. IV presents how this approach can be used in NR to simultaneously leverage the strengths of therapists and robotic devices. Sec. V outlines examples of potential applications of these technologies in the context of conventional rehabilitation practice. Sec. VI discusses challenges associated with the implementation of these technologies. Finally, Sec. VII explores open questions and future research directions.
@@ III Taxonomy and Overview of RHHI in NR
In RHHI, one or more robotic devices are used to render specific interaction dynamics between individuals [Ganesh2014, Takagi2018HapticInteraction, kuccuktabak2021human]. In NR, at least one individual is a patient physically connected to a robot, whose movements are influenced in real time by another user, such as a therapist, another patient, or even a family member. Together, the individuals engage in coordinated movements aimed at promoting motor learning for the patient(s) and facilitating the transfer of these skills to everyday functional activities. To better analyze RHHI scenarios in NR, the following section introduces key characterizations.
@@ III-A Interaction Configurations
Task-Space interactions, where users exchange forces through the end-effector of a robot (Fig. 1.A and Fig. 1.G), without specific postural corrections. Single-joint interactions, where the two users interact by moving a single joint (e.g., in Fig. 1.B, the wrist). This interaction is mediated using a single-DoF robots and allows joint-specific training. Multi-joint interactions allows coordination across multiple DoFs (Fig. 1.C and Fig. 1.E). This type of interaction is typically achieved with exoskeletons. Cross-limb interactions, in which the involved parts of the bodies of the two persons are not the same, such as the upper body of one user is connected to the lower body of another (Fig. 1.G). Guided interactions, where one person interacts with a robot and another person can influence the robot’s behavior, for example, by teaching a kinematic or force profile while directly interacting with the robot (Fig. 1.D and discussed in Box. VI-A).
RHHI can accommodate various interaction configurations, including bidirectional and unidirectional interaction. The interaction is defined as bidirectional if both users perceive interaction forces due to the interaction between them. This configuration can be implemented, for example, by either using two robotic devices that both render interaction forces to each user (Fig. 1.E) or using a transparent robotic device that directly transmits forces applied by one user to a second user (Fig. 1.D). Unidirectional interaction refers to a setup in which only one user perceives forces through a robotic device, while the device’s motion is teleoperated or influenced by another user. This can be achieved in real time (Fig. 1.E,F) or via prior demonstrations and replication using a single device (Fig. 1.D), either teleoperated or influenced by another user, or with two similar devices, where only one renders forces.
@@ III-B Interaction Medium
In RHHI, the interaction medium refers to the virtual connection implemented between two users through robotic devices. This connection can be designed to render specific mechanical behaviors, such as a virtual spring-damper model between joint positions, that enables users to feel each other’s movements through adjustable mechanical impedance levels [Takagi2018HapticInteraction, Che2016TwoCoupling, Ganesh2014, Tanaka2019ATask, Piovesan2013HapticTele-interaction]. A rigid virtual spring constant results in closely coupled motion [Takagi2018HapticInteraction, short2023haptic]. In contrast, a compliant, softer virtual spring allows more independent movement while still facilitating the exchange of forces. An asymmetric interaction medium can also be implemented, such as the case where one user experiences a rigid connection while the other experiences a soft one [noccaro2025robot], enabling customized feedback and workload distribution. It is important to note that the resulting behavior is not solely determined by the virtual model, but also by the human biomechanics, such as the stiffness of the involved joints, residual strength, and active participation of each user. Indeed, individuals can modulate their joint stiffness to either resist or follow the forces transmitted through the interaction [borner2023physically].
Figure 1: Examples of Robot-mediated physical Human-Human Interaction (RHHI) in neurorehabilitation (NR): (A) Two users (typically a patient and a partner such as a therapist, another patient, or a family member, Partner Characteristics) interact through two robots. The Task definition influences the User strategies and their performance. (B) Multiple patients can perform interactive tasks by virtually connecting multiple devices. (C) Therapist and patient interact via two upper-limb exoskeletons. They can assume different roles during the interaction (Interaction Scenario). (D) Robotic devices can be used for learning by demonstration strategies. (E-F) The interaction medium between the two users is modeled using a virtual spring-damper system characterized by therapist and patient stiffness ($K_{T},K_{P}$) and damping ($D_{T},D_{P}$) during bi-directional (E) and uni-directional (F) interaction. The interaction between users can be in joint-space (E) or in task-space (G).
@@ III-C Partner Characteristics
Partner characteristics refer to individual and relational factors that influence how users behave during RHHI. These include the relative skill levels of each user, as well as their social relationship (e.g., strangers, friends, family members, peers, or in a patient-therapist dynamic) [Che2016TwoCoupling, AvilaMireles2017SkillInteraction, Ganesh2014, Kager2019TheTask, Takagi2019IndividualsGoal, Mace2017BalancingTraining, B2018HapticField, Gorsic2017CompetitiveIntensity, Gorsic2017AGame]. Such characteristics can affect muscular effort, response strategies, task engagement, and the roles that emerge during the interaction. For example, a more skilled or confident partner may naturally take on a guiding role, while a less experienced partner may assume a more passive role. In the case of peer-based interactions, e.g., between two patients, it is particularly valuable to enable the scaling of targets, task difficulty, and interaction roles to address the individualized therapeutic objectives of each participant.
Another important dimension of the framework is the number of users participating in RHHI. Most studies focus on dyadic interactions, in which two individuals are physically connected through one or two robotic devices. However, multi-user configurations have also been explored [Takagi2019IndividualsGoal]. While these setups present additional challenges in terms of control complexity and task coordination, they offer unique opportunities for social engagement, scalability, and group-based rehabilitation.
@@ III-D Interaction Scenario
The interaction scenario defines the way the task is structured and how roles are assigned between participants [Jarrasse2012]. Common scenarios include collaboration, cooperation, and competition. These modes of interaction can significantly shape user behavior, influencing engagement, learning strategies, and the physical dynamics of interaction. In collaborative scenarios, both users share the same task goal and work together without predefined roles [Che2016TwoCoupling]. For example, a patient may perform a task jointly with a therapist, another patient, or a family member, sharing the effort required to complete movements that might otherwise be too difficult alone. In cooperative scenarios, users still share a goal, but with distinct roles such as teacher/student [Aoyagi2007, Banala2007, Kim2010] or leader/follower [Chackochan2019IncompleteAction]. A therapist, for instance, may lead the interaction by providing guidance while the patient adopts a more passive role. Alternatively, in a home scenario, a family member could guide a patient with the help of a robot [Baur2018RobotSupportedTraining]. Finally, in a clinical setting, a less severely impaired patient could potentially guide and support a more severely impaired patient or could take on a more difficult role in a shared task [GorsicRehab2020]. In contrast, competitive scenarios involve opposing goals, often implemented through game-based tasks [Gorsic2017AGame]. Competitive strategies can be employed to encourage patients’ active engagement and augment the challenge of a given task, though this would require careful balancing of exercise difficulty to ensure that, for example, a less impaired patient does not immediately overwhelm a more severely impaired patient [Gorsic2017AGame].
@@ III-E Task Performance and Motor Learning
Task performance in RHHI quantifies how users perform the assigned motor task while physically connected. Performance is typically measured through metrics such as accuracy, timing, smoothness, or synchronization. It is influenced by the design of the interaction medium, the characteristics of the users, and the selected interaction scenario. Physical feedback exchanged between users has been shown to improve accuracy, smoothness, coordination and overall task efficiency [Reed2006HapticMachines, Mojtahedi2017OnDyads, Takagi2018HapticInteraction, ivanova2022interaction].
Motor learning refers to persistent improvements in motor performance due to the RHHI training, after the physical connection is removed. This lasting change represents the core therapeutic effect in NR, reflecting the patient’s ability to retain and apply improved motor strategies independently. This is often assessed by comparing individual performance (i.e., without the physical interaction) before and after training. RHHI has been shown to support motor learning through various mechanisms, including increased motivation and engagement [Gorsic2017CompetitiveIntensity, GorsicRehab2020, GajadharDoubleFun], the opportunity to estimate and internalize a partner’s movement strategy [Takagi2017, Takagi2018HapticInteraction, Takagi2019IndividualsGoal] and role specialization that emerges during interaction [Reed2006HapticMachines, Mojtahedi2017OnDyads]. These benefits make RHHI particularly promising for NR, where one of the users aims to regain lost motor functions through repeated, adaptive practice.